OpenStudy (anonymous):

e^2x+1=20 solve for x An explanation would be greatly appreciated

OpenStudy (anonymous):

e^2x+1=20 (take 1 to other side) e^2x=19 (multiply both sides by ln to bring 2x exponent down infront of e) ln(e^2x)=ln(19) (ln and e cancel each other out always) 2x = ln19 x=ln19/2 x=1.4722

OpenStudy (anonymous):

That's what I found as well, I did manage to figure it out, but having somebody else confirm it is nice, thanks for the help!! :)

OpenStudy (anonymous):

no problem :)